# A dimensional mass transference principle from balls to open sets and applications to dynamical Diophantine approximation

### v1

@@ 3.2. Hausdorff content bound of $\bigcup_{k=\ell}^{\infty}E_{k}\cap B$
$\mathcal{H}_{\infty}^{s}(E_{n})\gg r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}.$
For any $n\geq 1$, by Frostman’s lemma, there exists a probability measure $\lambda_{n}$ supported on $E_{n}$ such that
(3.13) $\lambda_{n}(B(x,r))\ll\frac{r^{s}}{r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}}.$
Since $E_{n}\subset B(x_{n},r_{n})$, it follows that
$s\leq\operatorname{\overline{\dim}_{H}}\nu\leq\delta.$
@@
Next, we estimate the $\eta$-measure of arbitrarily balls, which will allow us to apply the mass distribution principle and conclude the desired lower bound on the Hausdorff content. Suppose that $r>0$ and
@@
Case 1: $r>|B|$. Since $\eta$ is a probability measure supported on a subset of $B$, the measure of any ball with radius larger than $|B|$ is trivially bounded by 1. Using the $\delta$-Ahlfors regularity of $\mu$, we have
@@
(3.17) $\eta(B(x,r))\leq\sum_{\begin{subarray}{c}F_{i}\in\mathcal{F}_{B}\\ F_{i}\subset B(x,2r)\end{subarray}}\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\ll\frac{\mu(B(x,2r))}{\mu(B)}\ll\frac{r^{\delta}}{\mu(B)}\ll\frac{r^{s}}{\mu(B)}.$
@@
(3.18) $\eta(B(x,r))\leq\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\ll\frac{|F_{i}|^{\delta}}{\mu(B)}\leq\frac{(16CNr)^{\delta}}{\mu(B)}\ll\frac{r^{s}}{\mu(B)},$
@@
$\begin{split}\eta(B(x,r))&\leq\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ B(x,r)\cap E_{n_{z}}\neq\emptyset\end{subarray}}\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\cdot\frac{\nu(L_{n_{z}})}{\sum_{L_{n_{z}}\in\mathcal{B}(F_{i})}\nu(L_{n_{z}})}\\ &\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ B(x,r)\cap E_{n_{z}}\neq\emptyset\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}.\end{split}$
@@
where the last inclusion uses the fact that $x\in E_{n_{w}}\subset B(w,5r_{n_{w}})$ (see (3.14)). It follows that
$\begin{split}\eta(B(x,r))&\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ B(x,r)\cap E_{n_{z}}\neq\emptyset\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}.\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ L_{n_{z}}\subset B(w,16r)\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}\\ &\leq\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(B(w,16r)\cap F_{i})}{\nu(F_{i})}\end{split}$
Note that $w\in E_{\nu,F_{i}}^{N,\varepsilon}$. Since $r<|F_{i}|/(16CN)$ (equivalently $16r<|F_{i}|/(CN)$), by the definition of $E_{\nu,F_{i}}^{N,\varepsilon}$,
$\frac{\mu(B(w,16r)\cap F_{i})}{\nu(F_{i})}\ll\bigg{(}\frac{16r}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}\ll\frac{r^{s-\varepsilon}}{|F_{i}|^{\delta}},$
where we use $s\leq\operatorname{\overline{\dim}_{H}}\nu\leq\delta$ in the last step. Putting all these together, we conclude that
(3.20) $\eta(B(x,r))\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(B(w,16r)\cap F_{i})}{\nu(F_{i})}\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{r^{s-\varepsilon}}{|F_{i}|^{\delta}}\ll\frac{r^{s-\varepsilon}}{\mu(B)}.$
Case 4: $0<r<r_{n_{w}}$. In this scale, due to the separation of the sets $E_{n_{z}}$ (see (3.12)), the ball $B(x,r)$ can intersect only the single set $E_{n_{w}}$ containing $x$. By the Frostman-type property for $\lambda_{n_{w}}$ (see (3.13)),
$\eta(B(x,r))$ $\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{w}})}{\nu(F_{i})}\cdot\lambda_{n_{w}}(B(x,r))\ll\frac{|F_{i}|^{\delta}}{\mu(B)}\cdot\bigg{(}\frac{r_{n_{w}}}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}\cdot\frac{r^{s}}{r_{n_{w}}^{\operatorname{\overline{\dim}_{H}}\nu}}$
(3.21) $\ll\frac{r_{n_{w}}^{-\varepsilon}r^{s}}{\mu(B)}\leq\frac{r^{s-\varepsilon}}{\mu(B)}.$
@@
4. Application to $\beta$-transformation
@@ 4.1. Definition and some basic properties
Denote by $\Sigma_{\beta}^{n}$ the collection of all admissible sequences of length $n$.
@@
where $\Lambda_{\beta}^{n}$ denotes the set of $n$th level full cylinders. In particular, the $\limsup$ set defined by all full cylinders has full Lebesgue measure.
@@ 4.2. Pressure function
where for each admissible word $\bm{\epsilon}_{n}$, the point $y$ is any element in the corresponding cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$, and $S_{n}\phi(y)$ denotes the ergodic sum $\sum_{k=0}^{n-1}\phi(T_{\beta}^{k}y)$. The existence of the limit in (4.1) follows from the subadditivity:
@@
and the limit does not depend on the choice of $y$ by the continuity of $g$.
It follows directly from the definition that the pressure function is continuous with respect to $\phi$.
@@
$|P(\phi,T_{\beta})-P(\varphi,T_{\beta})|\leq\sup_{x\in[0,1]}|\phi(x)-\varphi(x)|.$
Consequently, if $\phi$ is positive, then there exists $0<s=s(\phi)<1$ such that
@@
    For any cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ of level $n$, the measure $\nu_{\phi}$ satisfies the Gibbs upper bound:
@@
It is important to note that in condition (3) of Theorem 1.4, one needs to compare the Hausdorff content $\mathcal{H}_{\infty}^{s}(E_{n})$ with $r_{n}^{\operatorname{\overline{\dim}_{H}}\nu}$. This comparison relies on understanding the Hausdorff dimension of the Gibbs measure $\nu_{\phi}$. The author believes that the following result has been established elsewhere; however, since no suitable reference could be found, we include the proof here.
@@
By Birkhoff’s ergodic theorem, for $\nu_{\phi}$ almost all $x$,
@@
For any $n\geq 1$, denote by $I_{n,\beta}(x)$ the cylinder of level $n$ that contains $x$. For any $x$ for which (4.2) holds, by the definition of local dimension,
$\begin{split}\underline{D}(\nu_{\phi},x)&=\liminf_{r\to 0}\frac{\log\nu_{\phi}(B(x,r))}{\log r}\leq\liminf_{n\to\infty}\frac{\log\nu_{\phi}(I_{n,\beta}(x))}{\log|I_{n,\beta}(x)|}\\ &\leq\liminf_{\begin{subarray}{c}n\to\infty\\ I_{n,\beta}(x)\text{ if full}\end{subarray}}\frac{\log\nu_{\phi}(I_{n,\beta}(x))}{\log|I_{n,\beta}(x)|},\end{split}$
where, in the last inequality, we use the fact that for any $x\in[0,1)$, there exists infinitely many $n$ such that $I_{n,\beta}(x)$ is full (see Lemma 4.4). For any full cylinder $I_{n,\beta}(x)$, by the Gibbs property,
$\nu_{\phi}(I_{n,\beta}(x))\asymp e^{S_{n}\phi(x)-nP(\phi,T_{\beta})}=e^{n(\int\phi\,\mathrm{d}\nu_{\phi}-P(\phi,T_{\beta})+o(1))},$
@@
$\begin{split}\underline{D}(\nu_{\phi},x)&\leq\liminf_{\begin{subarray}{c}n\to\infty\\ I_{n,\beta}(x)\text{ if full}\end{subarray}}\frac{\int\phi\,\mathrm{d}\nu_{\phi}-P(\phi,T_{\beta})+o(1)}{-\log\beta}=\frac{P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}}{\log\beta}\\ &=\frac{h_{\nu_{\phi}}}{\log\beta},\end{split}$
where the equality follows from the variational principle. Since this holds for $\nu_{\phi}$ almost all $x$, we have
@@
It follows from the definition that
@@
Fix a Borel set $E$ with $\nu_{\phi}(E)>0$. Let $\varepsilon>0$. Then, there exists an integer $N=N(\varepsilon)$ such that the set $A_{N}$ of $x$ for which
@@
We stress that the reverse implication may not be true. Let $I_{n,\beta}(\bm{\epsilon}_{n})$ be a cylinder with positive $\nu_{\phi}$ measure and let $x=x(\bm{\epsilon}_{n})$ be such that (4.5) holds. Let $m\geq n$ be the unique integer satisfy
@@
It follows that $I_{n,\beta}(\bm{\epsilon}_{n})=I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n})$, where $0^{m-n}$ denotes the word consisting of mnm-n consecutive zeros. Since $x\in I_{n,\beta}(\bm{\epsilon}_{n})=I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n})$, it follows from Theorem 4.7 that
$\begin{split}\lambda(I_{n,\beta}(\bm{\epsilon}_{n}))&\leq\nu_{\phi}(I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n}))\ll e^{S_{m}\phi(x)-mP(\phi,T_{\beta})}\\ &\leq e^{m\int\phi\,\mathrm{d}\nu_{\phi}+m\varepsilon-mP(\phi,T_{\beta})}=\beta^{-m(P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}-\varepsilon)/\log\beta}\\ &\asymp|I_{n,\beta}(\bm{\epsilon}_{n})|^{(P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}-\varepsilon)/\log\beta}.\end{split}$
@@
$\operatorname{\dim_{H}}(A_{N}\cap E)\geq\frac{P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}-\varepsilon}{\log\beta}=\frac{h_{\nu_{\phi}}-\varepsilon}{\log\beta}.$
By the arbitrariness of $\varepsilon$, the claim (4.4) follows immediately. ∎

### v2

@@ 3.2. Hausdorff content bound of $\bigcup_{k=\ell}^{\infty}E_{k}\cap B$
(3.15) $\mathcal{H}_{\infty}^{s}(E_{n})\gg r_{n}^{\operatorname{\overline{\dim}_{H}}\nu},$
where the implied constant is independent of $n$. For any $n\geq 1$, by Frostman’s lemma, there exists a probability measure $\lambda_{n}$ supported on $E_{n}$ such that
[…]
where, again, the implied constant is independent of $n$.
Since $E_{n}\subset B(x_{n},r_{n})$, this immediately implies
(3.17) $s\leq\operatorname{\overline{\dim}_{H}}\nu\leq\delta.$
Indeed, suppose for contradiction that $s>\operatorname{\overline{\dim}_{H}}\nu$. Then,
[…]
This is impossible, since $|E_{n}|\to 0$ as $n\to\infty$ and the implied constants are independent of $n$. Hence (3.17) must hold.
@@
Remark 9.
The construction of the measure $\eta$ is fairly standard:
  1. (1)
    the sum of the $\mu$-measures of the sets $F_{i}\in\mathcal{F}_{B}$ is comparable to that of $B$ (see (3.12)). We therefore assign to each $F_{i}$ a weight equal to the proportion of its $\mu$-measure in the total $\mu$-measure, that is,
    $\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})};$
  2. (2)
    for each $F_{i}\in\mathcal{F}_{B}$, the sum of the $\nu$-measure of the sets $L_{n_{z}}\in F_{i}$ is comparable to that of $F_{i}$ (see (3.13)). We therefore assign to each $L_{n_{z}}\in F_{i}$ a weight
    $\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\cdot\frac{\nu(L_{n_{z}})}{\sum_{L_{n_{z}}\in\mathcal{B}(F_{i})}\nu(L_{n_{z}})};$
  3. (3)
    since $L_{n_{z}}$ has only one descendant $E_{n_{z}}$, we assign the entire weight of $L_{n_{z}}$ to $E_{n_{z}}$. This together with a natrual measure $\lambda_{n_{z}}$ supported on $E_{n_{z}}$ gives the definition of $\eta$.
Next, we estimate the $\eta$-measure of arbitrary balls, which will allow us to apply the mass distribution principle and conclude the desired lower bound on the Hausdorff content (see (3.1)). Suppose that $r>0$ and
@@
Case 1: $r>|B|$. Since $\eta$ is a probability measure supported on a subset of $B$, the measure of any ball with radius larger than $|B|$ is trivially bounded by 1. Using the $\delta$-Ahlfors regularity of $\mu$ and $s\leq\delta$ (see (3.17)), we have
@@
[…]
@@
[…]
@@
[…]
@@
where the last inclusion uses the fact that $x\in E_{n_{w}}\subset B(w,5r_{n_{w}})$ (see (3.18)). It then follows that
$\begin{split}\eta(B(x,r))&\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ B(x,r)\cap E_{n_{z}}\neq\emptyset\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}.\ll\sum_{\begin{subarray}{c}L_{n_{z}}\in\mathcal{B}(F_{i})\\ L_{n_{z}}\subset B(w,16r)\end{subarray}}\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(L_{n_{z}})}{\nu(F_{i})}\\ &\leq\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(B(w,16r)\cap F_{i})}{\nu(F_{i})}.\end{split}$
Note that $w\in E_{\nu,F_{i}}^{N,\varepsilon}$ by our construction (see Lemma 3.1 (2) and (3.18)). Since $r<|F_{i}|/(16CN)$ (equivalently $16r<|F_{i}|/(CN)$), by the definition of $E_{\nu,F_{i}}^{N,\varepsilon}$,
[…]
Putting these estimates together, we conclude that
(3.25) $\eta(B(x,r))\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{\nu(B(w,16r)\cap F_{i})}{\nu(F_{i})}\ll\frac{\mu(F_{i})}{\mu(B)}\cdot\frac{r^{s-\varepsilon}}{|F_{i}|^{\delta}}\asymp\frac{r^{s-\varepsilon}}{\mu(B)}.$
Case 4: $0<r<r_{n_{w}}$. In this scale, due to the separation of the sets $E_{n_{z}}$ (see (3.14)), the ball $B(x,r)$ can intersect only the single set $E_{n_{w}}$ containing $x$. Note that the discussion of Case 3 implies that for $r=r_{n_{w}}$ (see (3.24)),
$\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\cdot\frac{\nu(L_{n_{z}})}{\sum_{L_{n_{z}}\in\mathcal{B}(F_{i})}\nu(L_{n_{z}})}\ll\frac{r_{n_{w}}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}}{\mu(B)}.$
By the Frostman-type property for $\lambda_{n_{w}}$ (see (3.16)),
$\eta(B(x,r))$ $\leq\frac{\mu(F_{i})}{\sum_{F_{i}\subset\mathcal{F}_{B}}\mu(F_{i})}\cdot\frac{\nu(L_{n_{z}})}{\sum_{L_{n_{z}}\in\mathcal{B}(F_{i})}\nu(L_{n_{z}})}\cdot\lambda_{n_{w}}(B(x,r))$
[…]
(3.26) $=\frac{r_{n_{w}}^{-\varepsilon}r^{s}}{\mu(B)}\leq\frac{r^{s-\varepsilon}}{\mu(B)},$
where the last inequality follows from $0<r<r_{n_{w}}$.
@@
4. Application to $\beta$-transformations
@@ 4.1. Definition and some basic properties
Denote by $\Sigma_{\beta}^{n}$ the collection of all admissible sequences of length $n$. The following result of Rényi [40] implies that the dynamical system $([0,1],T_{\beta})$ admits $\log\beta$ as its topological entropy.
Lemma 4.2 ([40, (4.9) and (4.10)]).
Let $\beta>1$. For any $n\geq 1$,
$\beta^{n}\leq\#\Sigma_{\beta}^{n}\leq\frac{\beta^{n+1}}{\beta-1},$
where #\# denotes the cardinality of a finite set.
@@
The key property of full sequences is that the concatenation of any two full sequences is again full.
Proposition 4.5 ([19, Lemma 3.2]).
An $n$th level cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ is full if and only if, for any $\beta$-admissible sequence $\bm{\epsilon}^{\prime}_{m}\in\Sigma_{\beta}^{m}$ with $m\geq 1$, the concatenation $\bm{\epsilon}_{n}\bm{\epsilon}_{m}^{\prime}$ is still $\beta$-admissible. Moreover,
$|I_{n+m,\beta}(\bm{\epsilon}_{n}\bm{\epsilon}_{m}^{\prime})|=|I_{n,\beta}(\bm{\epsilon}_{n})|\cdot|I_{m,\beta}(\bm{\epsilon}_{m}^{\prime})|.$
A corollary in [19] establishes a certain relationship between full and non-full cylinders. While the statement given below is not stated exactly as in their original work, it can nonetheless be rigorously justified using the same method.
Lemma 4.6 ([19, Corollary 3.3 and (3.4)]).
Let $n\in\mathbb{N}$ and $\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}$. Let $m\geq n$ be the unique integer satisfying
$\beta^{-m-1}<|I_{n,\beta}(\bm{\epsilon}_{n})|\leq\beta^{-m}.$
Then, $I_{n,\beta}(\bm{\epsilon}_{n})=I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n})$ and $I_{m+1}(\bm{\epsilon}_{n},0^{m-n+1})\subset I_{n,\beta}(\bm{\epsilon}_{n})$ is full, where $0^{k}$ denotes the word consisting of $k$ consecutive zeros.
@@
where $\Lambda_{\beta}^{n}$ denotes the set of full sequences of length $n$. In particular, the $\limsup$ set defined by all full cylinders has full Lebesgue measure.
@@ 4.2. Pressure function
where for each admissible word $\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}$, the point $y$ is any element in the corresponding cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$, and $S_{n}\phi(y)$ denotes the ergodic sum $\sum_{k=0}^{n-1}\phi(T_{\beta}^{k}y)$. The existence of the limit in (4.1) follows from the subadditivity:
@@
and the limit does not depend on the choice of $y$ by the continuity of $\phi$.
It follows directly from the definition that the pressure function is continuous with respect to $\phi$. In the absence of a suitable reference, we provide a proof for the sake of completeness.
@@
$|P(\phi,T_{\beta})-P(\varphi,T_{\beta})|\leq\|\phi-\varphi\|_{\infty},$
where $\|\cdot\|_{\infty}$ denotes the supremum norm of a function in $C^{0}([0,1])$. Consequently, if $\phi$ is positive, then there exists $0<s=s(\phi)<1$ such that
@@
Proof.
By the definition of $P(\phi,T_{\beta})$ (see (4.1)), for each $n\in\mathbb{N}$, we have
$\begin{split}\frac{1}{n}\log\sum_{\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}}e^{S_{n}\phi(y)}&=\frac{1}{n}\log\sum_{\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}}e^{S_{n}\varphi(y)+S_{n}\phi(y)-S_{n}\varphi(y)}\\ &\leq\frac{1}{n}\log\sum_{\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}}e^{S_{n}\varphi(y)+n\|\phi-\varphi\|_{\infty}}\\ &=\bigg(\frac{1}{n}\log\sum_{\bm{\epsilon}_{n}\in\Sigma_{\beta}^{n}}e^{S_{n}\varphi(y)}\bigg)+\|\phi-\varphi\|_{\infty}.\end{split}$
Letting $n\to\infty$, it follows that
$P(\phi,T_{\beta})-P(\varphi,T_{\beta})\leq\|\phi-\varphi\|_{\infty}.$
Interchanging the roles of $\phi$ and $\varphi$ immediately yields the first point of the proposition.
By the above conclusion, the function $s\mapsto P(-s(\phi+\log\beta),T_{\beta})$ is continuous in $s$. By the intermediate value theorem, to conclude the second point of the proposition, it suffices to prove that $P(-s(\phi+\log\beta),T_{\beta})$ takes values of opposite signs at $s=0$ and $s=1$. For $s=0$, we have
[…]
For $s=1$, since $\phi$ is positive, we have
[…]
which is what we want. ∎
@@
    For any cylinder $I_{n,\beta}(\bm{\epsilon}_{n})$ of level $n$, the measure $\nu_{\phi}$ satisfies the upper Gibbs property:
@@
    Here, the implied constants do not depend on the particular choice of $x\in I_{n,\beta}(\bm{\epsilon}_{n})$.
It is important to note that in condition (3) of Theorem 1.4, one needs to compare the Hausdorff content $\mathcal{H}_{\infty}^{s}(E_{n})$ with $r_{n}^{\operatorname{\overline{\dim}_{H}}\nu_{\phi}}$. This comparison relies on understanding the Hausdorff dimension of the Gibbs measure $\nu_{\phi}$. The author believes that the following result has been established elsewhere; however, since no suitable reference could be found, we include the proof here.
@@
For any $n\geq 1$, denote by $I_{n,\beta}(x)$ the cylinder of level $n$ that contains $x$. Obviously,
$I_{n,\beta}(x)\subset B(x,|I_{n,\beta}(x)|),$
and so
$\frac{\log\nu_{\phi}(B(x,|I_{n,\beta}(x)|))}{\log|I_{n,\beta}(x)|}\leq\frac{\log\nu_{\phi}(I_{n,\beta}(x))}{\log|I_{n,\beta}(x)|}.$
Note that $|I_{n,\beta}(x)|$ goes to 0 as $n\to\infty$. By the definition of local dimension,
$\begin{split}\underline{D}(\nu_{\phi},x)&=\liminf_{r\to 0}\frac{\log\nu_{\phi}(B(x,r))}{\log r}\leq\liminf_{n\to\infty}\frac{\log\nu_{\phi}(B(x,|I_{n,\beta}(x)|))}{\log|I_{n,\beta}(x)|}\\ &\leq\liminf_{n\to\infty}\frac{\log\nu_{\phi}(I_{n,\beta}(x))}{\log|I_{n,\beta}(x)|}\leq\liminf_{\begin{subarray}{c}n\to\infty\\ I_{n,\beta}(x)\text{ is full}\end{subarray}}\frac{\log\nu_{\phi}(I_{n,\beta}(x))}{\log|I_{n,\beta}(x)|},\end{split}$
where, in the last inequality, we use the fact that for any $x\in[0,1)$, there exists infinitely many $n$ such that $I_{n,\beta}(x)$ is full (see Lemma 4.7).
By Birkhoff’s ergodic theorem, for $\nu_{\phi}$-almost all $x$,
@@
Let $x\in[0,1]$ be such that (4.2) holds. For any full cylinder $I_{n,\beta}(x)$, by the Gibbs property (see Theorem 4.10 (2)),
[…]
@@
$\begin{split}\underline{D}(\nu_{\phi},x)&\leq\liminf_{\begin{subarray}{c}n\to\infty\\ I_{n,\beta}(x)\text{ is full}\end{subarray}}\frac{n\big(\int\phi\,\mathrm{d}\nu_{\phi}-P(\phi,T_{\beta})+o(1)\big)}{\log|I_{n,\beta}(x)|}=\frac{P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}}{\log\beta}\\ &=\frac{h_{\nu_{\phi}}}{\log\beta},\end{split}$
where the equality follows from the variational principle (see Theorem 4.10 (1)). Since this holds for $\nu_{\phi}$-almost all $x$, we have
@@
It then follows from the definition of $\operatorname{\underline{\dim}_{H}}\nu_{\phi}$ that
@@
Fix a Borel set $E$ with $\nu_{\phi}(E)>0$. Let $\varepsilon>0$. Then, in view of (4.2), there exists an integer $N=N(\varepsilon)$ such that the set $A_{N}$ of $x$ for which
@@
We stress that the reverse implication may not be true. Let $I_{n,\beta}(\bm{\epsilon}_{n})$ be a cylinder with positive $\nu_{\phi}$-measure and let $x=x(\bm{\epsilon}_{n})$ be such that (4.5) holds. Let $m\geq n$ be the unique integer satisfying
@@
By Lemma 4.6, we have $I_{n,\beta}(\bm{\epsilon}_{n})=I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n})$. Since $x\in I_{n,\beta}(\bm{\epsilon}_{n})=I_{m,\beta}(\bm{\epsilon}_{n},0^{m-n})$, it follows from Theorem 4.10 (2) that
[…]
@@
$\operatorname{\dim_{H}}(A_{N}\cap E)\geq\frac{P(\phi,T_{\beta})-\int\phi\,\mathrm{d}\nu_{\phi}-\varepsilon}{\log\beta}=\frac{h_{\nu_{\phi}}-\varepsilon}{\log\beta},$
where the equality follows from Theorem 4.10 (1). By the arbitrariness of $\varepsilon$, the claim (4.4) follows immediately. ∎